OpenStudy (anonymous):

Divide the following polynomial (27x^2-9a)/(3)

zepdrix (zepdrix):

Ok so we don't have to do anything fancy like long division for this one. We just need to remember how fractions work. \[\large \frac{a+b}{c} \qquad=\qquad \frac{a}{c}+\frac{b}{c}\]

zepdrix (zepdrix):

So for our problem, we have,\[\large \frac{27x^2-9a}{3} \qquad=\qquad \frac{27x^2}{3}+\frac{-9a}{3}\]

zepdrix (zepdrix):

From there, just divide each coefficient by 3. 27/3 = ?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

take look on screen the answer should fit on grid so they give 9 to start

zepdrix (zepdrix):

you got 9 for the first coefficient? Ok good. The x^2 won't change, only the coefficient.\[\large 9x^2+\frac{-9a}{3}\] How bout the other one? -9/3=?

OpenStudy (anonymous):

thats -3

zepdrix (zepdrix):

Ok good. \[\large 9x^2-3a\] Having any trouble putting the symbols in correctly?

OpenStudy (anonymous):

no

zepdrix (zepdrix):

k cool

OpenStudy (anonymous):

(6x^3+11x^2-4x-4)/(3x-2) im stuck on this one too same thing

OpenStudy (anonymous):

2x+7 is the answer to (6x^3+11x^2-4x-4)/(3x-2)